OpenStudy (archtangent):

in VECTOR ANALYSIS if a cross product and dot product are in one equation, which operation goes first??? like (PEMDAS) Ex. A.BxC which operation should i do first???

ganeshie8 (ganeshie8):

you better do cross product first

ganeshie8 (ganeshie8):

doing dot product first makes no sense because the dot product yields a scalar and you cant cross it with vector

OpenStudy (fibonaccichick666):

I concur

ganeshie8 (ganeshie8):

you can assume below in equations \[A\cdot B\times C = A\cdot (B\times C)\]

OpenStudy (archtangent):

hmmm makes sense thank you sirs appreciate it :)) thumbs up

ganeshie8 (ganeshie8):

yw:)

OpenStudy (perl):

actually (A . B) x C does not make sense A . (BXC) makes sense

OpenStudy (perl):

A.B is a scalar, and you cant cross a vector with a scalar. but BXC is a vector , so A . (BXC) makes sense. so you dont need parentheses here

OpenStudy (fibonaccichick666):

yep, it is because of the limitations of a cross product that this phenomenon occurs